OpenStudy (anonymous):

Factor completely: 2x^3 + 4x^2 + 6x + 12

OpenStudy (callisto):

2x^3 + 4x^2 + 6x + 12 = 2( x^3 + 2x^2 + 3x +6) = 2[x^2 (x+2) + 3(x+2)] = 2(x^2 +3) (x+2)